Hilton Enterprises sells a product for $116 per unit. The variable cost is $79 per unit, while fixed costs are $184,815. Determine (A) the break-even point in sales units and (B) the break-even point if the selling price were increased to $124 per unit. a.
Answer #1

a.Contribution margin=Sales-Variable cost  

=(116-79)=$37 per unit

Breakeven=Fixed expenses/Contribution margin

=184815/37

=4995 units

b.Contribution margin=Sales-Variable cost  

=(124-79)=$45 per unit

Breakeven=Fixed expenses/Contribution margin

=184815/45

=4107 units